OpenStudy (anonymous):

A cell phone company offers a plan that includes free evening and weekend calling, as well as 100 weekd4 inihutes. The cost of this plan is $30 per month, plus $0.30 for each extra minute used that is not included in the plan. The cost for cell phone use according to this plan is C = 30 + 0.30m, where C is the total cost and m is the number of additional minutes used. (Taxes are already included in these costs and DO NOT need to be calculated or added) a) How much would it cost in a month if you used an extra 100 minutes in addition to the plan? EQUATION QUESTION

OpenStudy (anonymous):

by them saying extra 100 minutes i get that they are saying 100 over what was given so it would be m=100 and just plug and chug C = 30 + (.30)(100) C = 30 + 30 = 60